# Balanced Similarity with Auxiliary Prompts: Towards Alleviating Text-to-Image Retrieval Bias for CLIP in Zero-shot Learning

### v1

@@ 4.2 Auxiliary Prompt Generation
Most methods use CLIP to align the texts and images. The texts would contain at least one object. Since we cannot know what would meet in zero-shot learning settings, we select the object categories in the existing datasets as the contents in the auxiliary prompts, as the existing datasets contain most of the commonly seen objects. We have tested the object categories in COCO [27], CIFAR [19], and Caltch [19] and experimentally found that the categories combined from COCO and CIFAR obtained better results.
@@
Moreover, we find that in several tasks [51, 50], the query texts are sentences with objects, such as ”piece of broccoli at far right”. We experimentally discovered that the template length would also affect the performance of CLIP. Consequently, we use ChatGPT [10] to adjust the fixed templates with the same length of the input query text. The prompt for ChatGPT to adjust is defined as follows: Provide me with a template of the number of words is $X$ similar to “a photo of <>absent<>< >”. Note that, in such situation, the query text will not be added with the fixed template.
4.3 Hybrid Similarity
We drew inspiration from [16] to adopt a hybrid similarity that combines our BSAP and the original similarity of CLIP to obtain more robust retrieval results. The hybrid similarity formula is as follows:
@@ 4.3 Hybrid Similarity
where $\alpha$ is a constant parameter, $\mathbf{BS}_{m}$ is the final balanced similarity score, and $\mathbf{S}_{\text{ref}_{m}}$ represents the original similarity considering only the textual referring expression and all the image regions:
@@
Hybrid similarity incorporates both the original similarity derived from focused query text and our BSAP which can alleviate bias in text-to-image retrieval, thereby achieving more stable results.
5 Apply BSAP to different zero-shot models
5.1 Apply BSAP to Reclip in REC Task
Figure 5: Apply BSAP to Reclip.
@@ 5.1 Apply BSAP to Reclip in REC Task
As shown in Figure 5, the blue section represents the original core of the ReCLIP model. The orange section denotes the components that our BSAP (Balanced Similarity with Auxiliary Proplts) integrates into ReCLIP. The red boxes highlight the original similarity and the balanced similarity. In the figure, the parser divides the sentence into multiple segments. Each caption contains a core word. The number of divisions corresponds to the number of times CLIP is used, and accordingly, our BSAP is applied the same number of times. The hybrid similarity obtained by combining the original similarity and balanced similarity serves as the input to the Spatial Relation Resolver, replacing the original IPS results.
Figure 6: Apply BSAP to Zero-shot-RIS.
5.2 Apply BSAP to Zero-shot-RIS in RIS Task
@@
As shown in Figure 5, the blue section represents the original core of the Zero-shot-RIS model. The orange section denotes the components that our BSAP (Balanced Similarity with Auxiliary Proplts) integrates into Zero-shot-RIS. The red boxes highlight the original similarity and the balanced similarity.
6 Experiment
We conducted extensive experiments to demonstrate the capability of our BSAP. Three questions have been answered in this section:
@@ 6 Experiment
  • Q2: Can our BSAP mitigate the problem of text-to-image retrieval bias for CLIP?
@@
6.1 Experimental Settings
We first introduce the experimental settings. Two typical zero-shot learning tasks are selected as the test bed: REC and RIS. The datasets, evaluation metrics, and implementation details are given subsequently.
6.1.1 Datasets
@@ 6.1.2 Evaluation Metrics
The Zero-shot-RIS model for the RIS task uses the overall Intersection over Union (oIoU) and the mean Intersection over Union (mIoU). oIoU is calculated by dividing the total area of intersection by the total area of the union, where the total area is computed by aggregating over all examples. Additionally, mIoUs calculates the average IoU across all examples while taking the object sizes into consideration, as reported in [7, 18].
6.1.3 Implementation Details
The core $t_{\text{head }}$ are selected for auxiliary prompts including 80 classes of COCO  [27] and 100 classes of CIFAR [19]. “a photo of” is consistently used as the template. Regarding the template length, the Reclip model undergoes further analysis and processing of the text, leading to variable lengths. For the Zero-shot-RIS model, the corresponding input remains unprocessed. Templates with a length of 2 are chosen for RefCOCO and RefCOCO+ datasets, and a length of 5 for RefCOCOg dataset. The generated templates of different lengths are all generated based on ChatGPT [10]. As for the constant parameter $\alpha$, $\alpha=0.75$ is selected for REC task and $\alpha=0.5$ is selected for RIS task. Since it is a zero-shot task, the constant parameter $\alpha$ in different tasks is only tested on the val data set of RefCOCOg. See the appendix for details. Better parameter values may exist in other data sets. Our strategy and the enhanced models were operated in an environment with one RTX-2080 GPU, requiring no additional training.
6.2 Improvements of CLIP with BSAP (Q1)
6.2.1 Performance on zero-shot REC
Table 1: Accuracy on the RefCOCOg, RefCOCO+ and RefCOCO datasets in zero-shot REC task. The best zero-shot results in each column are in bold, and the previously best results are underlined. Reclip uses an ensemble of the RN50x16 and ViT-B/32 CLIP models. CPT-adapted is an adapted version of CPT-Blk. Supervised model UNINEXT-H [6]. All methods except UNITER-L use ground truth proposals from official annotations of MSCOCO as candidates.
@@ 6.2.1 Performance on zero-shot REC
BSAP(Balanced Similarity) 67.73 68.33 59.80 65.40 52.96 60.90 64.72 58.47
BSAP(Hybrid Similarity) 69.65 68.93 57.23 60.37 54.00 57.57 58.58 58.27
In Table 1, the performance of Reclip [39] with balanced similarity and hybrid similarity is presented in referring expression comprehension (REC), compared to CPT-adapted [47], GradCAM [36], Reclip [39] on RefCOCO, RefCOCO+, and RefCOCO datasets. To avoid the influence of accuracy variations in proposals from different detection models, Ground-truth proposals of MSCOCO are utilized as input following Reclip.
When only employing the balanced similarity, there is a significant improvement observed on the RefCOCO+ and RefCOCO datasets, with the highest boost reaching $12.73\%$. Hybrid similarity also consistently outperforms the original baseline methods across the board. The results demonstrate that our BSAP has the ability to improve the CLIP on REC.
6.2.2 Performance on Zero-shot RIS
Table 2: Comparison with Zero-shot RIS baseline methods on three standard benchmark datasets in zero-shot RIS task. All baseline methods use FreeSOLO as the mask proposal network. Best zero-shot results in each column are in bold, and the previously best results are underlined. FreeSOLO upper-bound is computed between the GT mask and the maximum overlapped FreeSOLO mask with the GT mask.
@@ 6.2.2 Performance on Zero-shot RIS
BSAP(Balanced Similarity) ViT-B/32 31.07 31.07 26.14 25.72 25.39 25.05 24.80 24.55
BSAP(Hybrid Similarity) ResNet-50 30.14 29.91 26.28 25.10 25.71 24.89 24.01 24.61
BSAP(Hybrid Similarity) ViT-B/32 31.45 31.31 26.43 25.59 25.89 25.16 24.89 24.70
@@
BSAP(Balanced Similarity) ViT-B/32 34.31 34.56 28.86 28.39 28.35 27.59 27.28 27.28
BSAP(Hybrid Similarity) ResNet-50 33.28 33.12 28.86 27.70 28.50 27.34 26.36 27.31
BSAP(Hybrid Similarity) ViT-B/32 34.52 34.48 28.67 27.80 28.30 27.34 27.02 27.10
@@
TSEG [38] ${}^{\dagger}$ 25.95 - 22.62 - - 23.41 - -
In Table 2, the performance of Zero-shot-RIS with our methods in RIS task is presented, employing both balanced similarity and hybrid similarity, along with other baselines on RefCOCO, RefCOCO+, and RefCOCO datasets in terms of oIoU and mIoU metrics. For a fair comparison, FreeSOLO [43] mask proposals are used to generate the final output masks for all methods, including Region token [23], Cropping, and Zero-shot-RIS [49].
The experimental results indicate that our proposed method outperforms Zero-shot-RIS and other baseline methods comprehensively. The balanced similarity in BSAP outperforms the previous methods across all datasets in terms of mIoU metrics, while the hybrid similarity in BSAP excels the previous methods in all metrics across all datasets. Although there is still a gap compared to fully supervised reference image segmentation methods [46], our approaches significantly improve performance compared to baselines with similar upper bounds. Additionally, Hybrid Similarity outperforms supervised method [38].
Our BSAP can be treated as a plug-in module to be easily applied in CLIP. The improvements in REC and RIS on two state-of-the-art CLIP-based methods also validate the generalizability of our BSAP. More technical details on how to apply our BSAP in the two tasks can be found in the supplemental materials.
Table 3: The balanced similarity enhances CLIP’s discriminative ability in REC task, observed in the RefCOCOg, RefCOCO+, and RefCOCO validation datasets. (a). The proportion of predicted object with the same category as the ground truth. (b). The proportion of cases where the predicted object and the ground truth object are the same when their categories same.
Dataset Reclip BSAP
@@
Dataset Reclip BSAP
@@
6.3 On mitigating text-to-image retrieval bias using BSAP (Q2)
As aforementioned analysis, the text-to-image retrieval bias of CLIP can be addressed by balancing the original similarity with auxiliary prompts. To demonstrate the affection more clearly, we use GT (Ground Truth) from MSCOCO instances as our proposal boxes and mask proposals. We adopt Reclip in REC as the test method. Table 3(b) presents the extent to which the text-query balanced similarity enhances the recognition accuracy of the CLIP model in the REC task. Specifically, in (a), the results are the accuracy of the query texts being correctly matched with the image regions that contain the same categories. In (b), the results are the accuracy that the query texts being correctly matched with the image regions that contain not only the same categories but also the correct attribute description.
The recognition ability between different object categories in (a) has improved by an average of $4\%$, and the fine-grained recognition ability between objects of the same category in (b) has improved by an average of $9\%$. The improvements demonstrate that our BSAP could mitigate CLIP’s text-to-image retrieval bias and enable CLIP to obtain better retrieval performance.
To provide a more intuitive illustration, we present the visual representations using Reclip and Zero-shot-RIS as shown in Figure 7. It can be seen that with BSAP, the CLIP-based methods can locate the correct image regions better.
6.4 Ablation Study (Q3)
Table 4: In the validation datasets of RefCOCOg, RefCOCO+, and RefCOCO, we studied the effect of different dataset categories in the content of our balanced similarity auxiliary prompts on accuracy for the REC task.
@@ 6.4 Ablation Study (Q3)
6.4.1 Effects of Contents Selection.
Due to the nature of zero-shot tasks, there are no referring descriptions in the dataset in advance, making it impossible to design our auxiliary prompts based on these references. For the core words of auxiliary prompts, 80 categories from the COCO dataset [27], 100 categories from the CIFAR dataset [19], and 101 categories from the Caltech dataset [20] are tested. These datasets are widely used, and their object categories contain commonly seen objects. The impact of different quantities of auxiliary prompts is studied on the results using these categories, as shown in Table 4. In the end, 180 categories from the COCO dataset and CIFAR dataset are chosen as the core words for our auxiliary prompts. The number of auxiliary prompts is also set to 180.
Table 5: In the validation datasets of RefCOCOg, RefCOCO+, and RefCOCO, we studied the effect of different template lengths in of our balanced similarity auxiliary prompts on mIoU for the RIS.
@@
6.4.2 Effects of Template Length.
Whether the textual information input into the CLIP model is incorporated into the template can indeed impact the final results. Since we are dealing with a zero-shot problem, our template selection is limited to considering the length of the text. Prompts for ChatGPT [10] are used to generate templates of different lengths, e.g., “a photo of” and “this is”. By generating ten templates for each length, fluctuation in the results is observed, as shown in Table 5. Finally, template lengths of 2, 2, and 5 are selected for RefCOCO, RefCOCO+, and RefCOCOg, respectively. The specific templates that were generated can be found in the appendix.
Table 6: In the validation datasets of RefCOCOg, RefCOCO+, and RefCOCO, we studied the effect of different balanced methods for balanced similarity on our best performance accuracy in the REC.
@@
6.4.3 The impact of normalization strategies.

### v2

@@ IV-B Auxiliary Prompts Generation
Most methods use CLIP to align the texts and images. The texts would contain at least one object. Since we cannot know what would meet in zero-shot learning settings, we select object categories in the existing datasets as contents in the auxiliary prompts, as existing datasets contain most of the commonly seen objects. We have tested the object categories in COCO [36], CIFAR [37], and Caltch [37] and experimentally found that categories combined from COCO and CIFAR obtained better results.
@@
Moreover, we find that in several tasks [38, 39], the query texts are sentences with objects, such as “piece of broccoli at far right”. We experimentally discovered that the template length would also affect the performance of CLIP. Consequently, we use ChatGPT [40] to adjust the fixed templates with the same length of the input query text. The prompt for ChatGPT to adjust is defined as follows: Provide me with a template of the number of words is $X$ similar to “a photo of <>absent<>< >”. Note that, in such a situation, the query text will not be added with the fixed template.
IV-C Hybrid Scores Using BSAP
We drew inspiration from [41] to adopt hybrid scores that combined our balanced scores using BSAP and original similarity scores of CLIP to obtain more robust retrieval results. The hybrid scores formula is as follows:
@@ IV-C Hybrid Scores Using BSAP
where $\alpha$ is a constant parameter, $\mathbf{BS}_{m}$ is the final balanced score, and $\mathbf{S}_{\text{ref}_{m}}$ represents the original scores considering only the textual referring expression and all the image regions:
@@
Hybrid scores incorporate both the original scores derived from focused query text and balanced scores using our auxiliary prompts which can alleviate hallucinations in text-to-image retrieval, thereby achieving more stable results.
V Examples of Applying BSAP to different zero-shot models
V-A Applying BSAP to Reclip in REC Task
Figure 3: Apply BSAP to Reclip. Blue sections represent the original core of ReCLIP. Orange sections denote the components that our BSAP and BSAP-H integrate into ReCLIP. Red boxes highlight original scores and our BSAP.
@@ V-A Applying BSAP to Reclip in REC Task
As shown in Fig. 3, the blue section represents the original core of the ReCLIP model. The orange section denotes the components that our BSAP and BSAP-H integrate into ReCLIP. The red boxes highlight the original similarity scores and our BSAP. In the figure, the parser divides the sentence into multiple segments. Each caption contains a core word. The number of divisions corresponds to the number of times CLIP is used, and accordingly, auxiliary prompts are applied the same number of times. The hybrid scores (BSAP-H) serve as the input to the Spatial Relation Resolver, replacing the original IPS results.
Figure 4: Apply BSAP to Zero-shot-RIS. Blue sections represent the original core of Zero-shot-RIS. Orange sections denote components that our BSAP and BSAP-H integrate into Zero-shot-RIS. Red boxes highlight original scores and BSAP.
V-B Apply BSAP to Zero-shot-RIS in RIS Task
@@
As shown in Fig. 4, the blue section represents the original core of the Zero-shot-RIS model. The orange section denotes the components that our BSAP and BSAP-H integrate into Zero-shot-RIS. The red boxes highlight the original similarity scores and BSAP.
VI Experiment
We conducted extensive experiments to demonstrate the capability of our BSAP. Four questions have been answered in this section:
@@ VI Experiment
  • Q2: Can our BSAP truly mitigate the text-to-image retrieval hallucinations for CLIP?
@@
  • Q4: Can our BSAP mitigate the hallucinations and improve the performance of other VLMs?
Additionally, we investigated whether BSAP could alleviate hallucinations in the image-to-text retrieval task.
VI-A Experimental Settings
We first introduce the experimental settings. Two typical zero-shot learning tasks are selected as the test beds: REC and RIS. The datasets, evaluation metrics, and implementation details are given subsequently.
VI-A1 Datasets
@@ VI-A2 Evaluation Metrics
The Zero-shot-RIS model for the RIS task uses the overall Intersection over Union (oIoU) and the mean Intersection over Union (mIoU). oIoU is calculated by dividing the total area of intersection by the total area of the union, where the total area is computed by aggregating over all examples. Additionally, mIoUs calculates the average IoU across all examples while taking object sizes into consideration, as reported in [42, 43].
VI-A3 Implementation Details
The core $t_{\text{head }}$ are selected for auxiliary prompts including 80 classes of COCO [36] and 100 classes of CIFAR [37]. “a photo of” is consistently used as the template. Regarding the template length, the Reclip model undergoes further analysis and processing of the text, leading to variable lengths. For the Zero-shot-RIS model, the corresponding input remains unprocessed. Templates with a length of 2 are chosen for RefCOCO and RefCOCO+ datasets, and a length of 5 for RefCOCOg dataset. Generated templates of different lengths are all generated based on ChatGPT [40]. As for the constant parameter $\alpha$, $\alpha=0.75$ is selected for the REC task and $\alpha=0.5$ is selected for the RIS task. Since it is a zero-shot task, the constant parameter $\alpha$ in different tasks is only tested on the val data set of RefCOCOg. Better parameter values may exist in other data sets. Our strategy and the enhanced models were operated in an environment with one RTX-2080 GPU, requiring no additional training.
VI-B Improvements of CLIP with BSAP (Q1)
VI-B1 Performance on zero-shot REC
TABLE I: Accuracy on RefCOCOg, RefCOCO+, and RefCOCO in zero-shot REC. The best results in each column are in bold, and previously best results are underlined. Reclip adopts RN50x16 and ViT-B/32 CLIP models. CPT-adapted is an adapted version of CPT-Blk. Supervised model UNINEXT-H [44]. All methods except UNITER-L adopt GT proposals from official annotations of MSCOCO as candidates. When using other VLMs, they all integrated into Reclip, replacing the original CLIP model. ALBEF uses the 14M version, while BLIP uses the ViT-L version.
@@ VI-B1 Performance on zero-shot REC
ReCLIP + BSAP 67.73 68.33 59.80 65.40 52.96 60.90 64.72 58.47
ReCLIP + BSAP-H 69.65 68.93 57.23 60.37 54.00 57.57 58.58 58.27
Other VLMs
ALBEF [45] 72.23 72.45 63.78 65.45 61.45 63.57 65.49 61.79
ALBEF [45] + BSAP 72.98 73.06 66.24 68.78 64.49 65.79 68.99 63.96
BLIP [46] 73.13 72.95 64.78 67.29 62.42 65.09 67.78 64.32
BLIP [46] + BSAP 74.54 74.14 67.67 68.78 64.55 65.79 69.45 65.94
In Table I, the performance of Reclip [10] with BSAP and BSAP-H is presented in referring expression comprehension (REC), compared to CPT-adapted [29], GradCAM [31], Reclip [10] on RefCOCO, RefCOCO+, and RefCOCO datasets. To avoid the influence of accuracy variations in proposals from different detection models, Ground-truth proposals of MSCOCO are utilized as inputs following Reclip.
When only employing the BSAP, there is a significant improvement observed on the RefCOCO+ and RefCOCO datasets, with the highest boost reaching $12.73\%$. BSAP-H also consistently outperforms the original baseline methods across the board. The results demonstrate that our BSAP has the ability to improve the CLIP on REC.
VI-B2 Performance on Zero-shot RIS
TABLE II: Comparison with Zero-shot RIS baseline methods on three standard benchmark datasets in zero-shot RIS task. All baseline methods use FreeSOLO as the mask proposal network. The best zero-shot results in each column are in bold, and the previously best results are underlined. FreeSOLO upper-bound is computed between the GT mask and the maximum overlapped FreeSOLO mask with the GT mask. When using other VLMs, they all integrated into Reclip, replacing the original CLIP model. ALBEF uses the 14M version, while BLIP adopts the ViT-L version.
@@ VI-B2 Performance on Zero-shot RIS
Zero-shot-RIS + BSAP ViT-B/32 31.07 31.07 26.14 25.72 25.39 25.05 24.80 24.55
Zero-shot-RIS + BSAP-H ResNet-50 30.14 29.91 26.28 25.10 25.71 24.89 24.01 24.61
Zero-shot-RIS + BSAP-H ViT-B/32 31.45 31.31 26.43 25.59 25.89 25.16 24.89 24.70
@@
Other LLMs
ALBEF(14M) [45] ViT-B/16 31.49 31.25 27.45 26.49 26.89 25.46 25.41 25.43
ALBEF(14M) + BSAP ViT-B/16 31.65 31.36 27.69 26.54 27.10 25.59 25.56 25.47
BLIP [46] ViT-L/16 31.54 31.29 27.49 26.78 26.87 25.56 25.74 25.56
BLIP + BSAP ViT-L/16 31.67 31.40 27.65 26.80 26.97 25.64 25.91 25.64
@@
Zero-shot-RIS + BSAP ViT-B/32 34.31 34.56 28.86 28.39 28.35 27.59 27.28 27.28
Zero-shot-RIS + BSAP-H ResNet-50 33.28 33.12 28.86 27.70 28.50 27.34 26.36 27.31
Zero-shot-RIS + BSAP-H ViT-B/32 34.52 34.48 28.67 27.80 28.30 27.34 27.02 27.10
@@
Other VLMs
ALBEF(14M) [45] ViT-B/16 35.14 35.27 29.21 30.14 29.79 28.78 27.49 27.79
ALBEF(14M) + BSAP ViT-B/16 36.65 35.56 29.78 30.56 30.45 29.30 28.20 28.24
BLIP [46] ViT-L/16 36.49 36.75 30.46 30.04 30.49 29.02 29.14 28.45
BLIP + BSAP ViT-L/16 37.03 37.25 30.99 30.63 30.97 29.61 29.46 29.00
@@
TSEG [49] ViT-S/16 25.95 - 22.62 - - 23.41 - -
In Table II, the performance of Zero-shot-RIS with our methods in the RIS task is presented, employing both BSAP and BSAP-H, along with other baselines on RefCOCO, RefCOCO+, and RefCOCO datasets in terms of oIoU and mIoU metrics. For a fair comparison, FreeSOLO [48] mask proposals are used to generate the final output masks for all methods, including Region token [32], Cropping, and Zero-shot-RIS [14].
The experimental results indicate that our method outperforms Zero-shot-RIS and other baseline methods comprehensively. The original model with BSAP outperforms previous methods across all datasets in terms of mIoU metrics, while the model with BSAP-H excels previous methods in all metrics across all datasets. Although there is still a gap compared to fully supervised reference image segmentation methods [47], our approaches significantly improve performance compared to baselines with similar upper bounds. Additionally, Zero-shot-RIS with BSAP-H outperforms supervised method [49].
Our BSAP can be regarded as a plug-in module to be easily applied in CLIP. The improvements in REC and RIS on two state-of-the-art CLIP-based methods also validate the generalizability of our BSAP.
TABLE III: BSAP enhances CLIP’s discriminative ability in the REC task, observed in the RefCOCOg, RefCOCO+, and RefCOCO validation datasets. The proportion of the predicted object with the same category as the ground truth.
Dataset Reclip Reclip + BSAP
@@
TABLE IV: BSAP enhances CLIP’s discriminative ability in the REC task, observed in the RefCOCOg, RefCOCO+, and RefCOCO validation datasets. The proportion of cases where the predicted object and the ground truth object are the same when their categories same.
Dataset Reclip Reclip + BSAP
@@
VI-C On Mitigating Hallucinations in Text-to-Image Retrieval Tasks Using BSAP (Q2)
As aforementioned analysis, the text-to-image retrieval hallucination of CLIP can be addressed by balancing the original scores with auxiliary prompts. To demonstrate the effectiveness more clearly, we use GT (Ground Truth) from MSCOCO instances as our proposal boxes and mask proposals. We adopt Reclip in REC as the test method. Table III and Table IV present the extent to which the text-query balanced similarity enhances the recognition accuracy of the CLIP model in the REC task. Specifically, in Table III, the results are the accuracy of the query texts being correctly matched with the image regions that contain the same categories. In Table IV, the results are the accuracy that the query texts are correctly matched with the image regions that contain not only the same categories but also the correct attribute description.
The recognition ability between different object categories in Table III has improved by an average of $4\%$, and the fine-grained recognition ability between objects of the same category in Table IV has improved by an average of $9\%$. The improvements demonstrate that our BSAP could mitigate CLIP’s text-to-image retrieval hallucination and enable CLIP to obtain better retrieval performance.
As shown in Fig. 5, to provide a more intuitive illustration, we present the visual representations using Reclip and Zero-shot-RIS. It can be seen that with BSAP, the CLIP-based methods can locate the correct image regions better.
VI-D Ablation Study (Q3)
TABLE V: In validation datasets of RefCOCOg, RefCOCO+, and RefCOCO, we studied the effect of different dataset categories in contents of our BSAP on accuracy for the REC.
@@ VI-D Ablation Study (Q3)
TABLE VI: In the validation datasets of RefCOCOg, RefCOCO+, and RefCOCO, we studied the effect of different template lengths in BSAP on mIoU for the RIS.
@@
TABLE VII: In the validation datasets of RefCOCOg, RefCOCO+, and RefCOCO, we studied the effect of different balanced methods for BSAP on our best performance accuracy in the REC.
@@
TABLE VIII: Apply BSAP to CLIP in the image-to-text retrieval task using the CIFAR-10 dataset.
Class CLIP CLIP + BSAP
airplane 75 91 $\uparrow$
automobile 74 76 $\uparrow$
bird 78 81 $\uparrow$
cat 58 59 $\uparrow$
deer 54 94 $\uparrow$
dog 51 79 $\uparrow$
frog 63 92 $\uparrow$
horse 92 87
ship 92 96 $\uparrow$
truck 83 92 $\uparrow$
VI-D1 Effects of Contents Selection
Due to the nature of zero-shot tasks, there are no referring descriptions in datasets in advance, making it impossible to design our auxiliary prompts based on these references. For core words of auxiliary prompts, 80 categories from the COCO dataset [36], 100 categories from the CIFAR dataset [37], and 101 categories from the Caltech dataset [50] are tested. These datasets are widely used, and their object categories contain commonly seen objects. The impact of different quantities of auxiliary prompts is studied on the results using these categories, as shown in Table V. Finally, 180 categories from COCO dataset and CIFAR dataset are chosen as core words for our auxiliary prompts. The number of auxiliary prompts is also 180.
VI-D2 Effects of Template Length
Whether the textual information input into the CLIP model is incorporated into the template can indeed impact final results. Since we are dealing with a zero-shot problem, our template selection is limited to considering the length of the text. Prompts for ChatGPT [40] are used to generate templates of different lengths, e.g., “a photo of” and “this is”. By generating ten templates for each length, fluctuation in the results is observed, as shown in Table VI. Finally, template lengths of 2, 2, and 5 are selected for RefCOCO, RefCOCO+, and RefCOCOg, respectively.
VI-D3 The impact of Normalization Strategies.